OpenStudy (tiff9702):

A recipe that makes 6 servings calls for 1 1/5 cups of juice. How many cups of juice are there per serving?

OpenStudy (zpupster):

use the fact that 1 1/5 (by (1x5) + 1) over 5) is equal to 6/5

OpenStudy (tiff9702):

What?

OpenStudy (anonymous):

So 6 servings is 1 1/5 cups so 6=1 1/5 1 1/5=6/5 so 6=6/5 divide both sides by 6 1 serving=1/5 cup

OpenStudy (anonymous):

i typed fast omg

OpenStudy (anonymous):

thx for the medal

OpenStudy (tiff9702):

Np